OpenStudy (bohotness):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 5) and (5, 1) C. (10, 6) and (5, 1) D. (10, 6) and (1, 5)

sammixboo (sammixboo):

Let's start with choice A

sammixboo (sammixboo):

OK so now working with choice A, let's start with the first coordinates. (10, 4), 10 = x, 4 = y, so plug it in into the equation given 4 = 10 - 4

sammixboo (sammixboo):

Is 4 = 10 - 4 true?

OpenStudy (bohotness):

i dont think yea

sammixboo (sammixboo):

So what is 10 - 4?

sammixboo (sammixboo):

It's 6, right? 4 = 10 - 4 4 = 6 4 does not equal 6, so therefor it isn't true. That also tells us that A isn't our answer

OpenStudy (bohotness):

10-6 so thats worng

sammixboo (sammixboo):

So now onto choice B and the first coordinate given. (10, 5) 10 = x, 5 = y 5 = 10 - 4 Is this true?

OpenStudy (bohotness):

no

sammixboo (sammixboo):

So B isn't our answer either. Onto C and the first coordinate given, (10, 6), 10 = x, and 6 = y Plug them in 6 = 10 - 4 Is this true?

OpenStudy (bohotness):

yes

sammixboo (sammixboo):

Right, so now lets see if this is our final answer by doing the second coordinate given, (5, 1) where 5 = x and 1 = y 1 = 5 - 4 Is this true?

OpenStudy (bohotness):

yes

sammixboo (sammixboo):

OK so that means C is our answers since both coordinates are true!

OpenStudy (bohotness):

:)

sammixboo (sammixboo):

If you need any more help, you know what to do =3

OpenStudy (bohotness):

can you help me with more lol please sammic:3

sammixboo (sammixboo):

Sure =) Sorry I was taking my dog outside. Make sure to post it on a diff post

OpenStudy (bohotness):

kk